At a local college, 21 of the male students are smokers and 189 are non-smokers. Of the female students, 90 are smokers and 210 are non-smokers. A male student and a female student from the college are randomly selected for a survey. What is the probabillty that both are smokers?
Answer #1

Result:

male

female

Total

smokers

21

90

111

non smokers

189

210

399

Total

210

300

510

P( smoker/male ) =21/210

P( smoker/female) = 90/300

P( both are smokers) = P( smoker/male )* P( smoker/female)

=(21/210)*(90/300)

=0.0300